the length of a rectangle is 8in. more than its width. The perimeter of the rectangle is 54 cm. What are the width and length of the rectangle
Answer by checkley77(12569)

L=W+8
2L+2W=54
2(W+8)+2W=54
2W+16+2W=54
4W=54-16
4W=38
W=38/4
W=9.5 IN. IS THE WIDTH.
L=9.5+8
L=17.5 IN.
PROOF:
2*17.5+2*9.5=54
35+19=54
54=54